OpenStudy (dtan5457):

QUick trig check...half angle tan 112.5 degrees

OpenStudy (dtan5457):

I got -1/sqrt 2

OpenStudy (dtan5457):

correct?

OpenStudy (dtan5457):

sin(a)/1+cos(a) \[\frac{ \frac{ -1 }{ \sqrt{2} } }{ 1+\frac{ 1 }{ \sqrt{2} } }\]

OpenStudy (dtan5457):

multiplied the 1 by sqrt2/sqrt2

OpenStudy (dtan5457):

added it up, crossed out the sqrt of 2

OpenStudy (dtan5457):

got -1/sqrt2 over 1

OpenStudy (dtan5457):

@TheSmartOne

OpenStudy (dtan5457):

@freckles @ksanka

OpenStudy (dtan5457):

@Nnesha @perl

OpenStudy (amorfide):

\[\tan(\frac{ B }{ 2})=\frac{ sinB }{ 1+cosB }\]

OpenStudy (amorfide):

\[\frac{ -\frac{ 1 }{ \sqrt{2} } }{1-\frac{ 1 }{ \sqrt{2} } }\]

OpenStudy (amorfide):

\[\frac{ -1 }{ \sqrt{2}-1 }\]

OpenStudy (dtan5457):

thats it?

OpenStudy (amorfide):

yeah that is exac t answer

OpenStudy (amorfide):

is that clear or do you need something explained?

OpenStudy (dtan5457):

is my friend also right? also, how can i check my answer is correct using a calculator?

OpenStudy (amorfide):

do tan112.5 on a calculator and match it to your answer

OpenStudy (dtan5457):

im not getting the same answer

OpenStudy (amorfide):

yes your friend is correct

OpenStudy (amorfide):

the answer you gave was wrong, I gave the right answer

OpenStudy (dtan5457):

no i mean checking on the calculator

OpenStudy (amorfide):

tan(112.5) in degrees gives you -2.414 the answer I gave you gives you -1-sqrt(2) which is the same answer

OpenStudy (dtan5457):

oh yeah, i got it now. thanks again.

OpenStudy (dtan5457):

i should stop studying for 3 tests at once

OpenStudy (dtan5457):

i made a dumb mistake here

OpenStudy (amorfide):

you weer on the right track just one small mistake

OpenStudy (dtan5457):

lol yeah, im gonna go work on this more letter. i'm gonna take one test at a time. thanks man

OpenStudy (dtan5457):

*later

OpenStudy (amorfide):

take care and good luck